# Crossed products by compact group actions with the weak tracial Rokhlin property

### v1

@@ 4. Simplicity of the fixed point algebra and the crossed product
Let $(e_{j,k})_{j,k=1,2,\dots,d_{\pi}}$ be the standard matrix units for $B(\mathcal{H}_{\pi})$. Set
$a=(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\sqrt{\varphi})(\pi^{*})=\sum_{j,k=1}^{d_{\pi}}e_{j,k}\otimes\sqrt{\varphi}((\pi_{k,j})^{*})\in B(\mathcal{H}_{\pi})\otimes\overline{r(A_{\infty,\alpha}\cap A^{\prime})r}.$
Since $\pi\pi^{*}=1$,
$a^{*}a=$ ${}(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\sqrt{\varphi})(\pi)(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\sqrt{\varphi})(\pi^{*})$
$=$ ${}(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\sqrt{\varphi})(\pi\pi^{*})(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\sqrt{\varphi})(1)$
$=$ ${}1_{B(\mathcal{H}_{\pi})}\otimes r.$
Since $C(G)$ is amenable, by the Choi–Effros lifting theorem, for $\delta$, $F=\{1_{A}\}$, and $S=\{\pi_{j,k},(\pi_{j,k})^{*}\colon j,k=1,2,\dots,d_{\pi}\}$, we obtain a contractive positive element $g\in A^{\alpha}$ such that $(1-g-\delta)_{+}\precsim_{A^{\alpha}}g$, and an $(S,F,\delta)$-approximately equivariant central strongly order zero completely positive contractive map $\psi\colon C(G)\to\overline{gAg}$ with $\psi(1)=g^{\frac{1}{2}}$ such that $\|g^{\frac{1}{2}}\psi(f_{1}f_{2})-\psi(f_{1})\psi(f_{2})\|<\delta$ for all $f_{1},f_{2}\in S$. As in Lemma 3.6, we will assume $\psi$ is equivariant. Set
@@
Since $\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\psi$ is equivariant and $\pi^{*}\in C(G)_{2}(\pi)$, a straightforward calculation shows that $c\in A_{2}(\pi)$. Indeed, for $g\in G$, since $\pi^{*}=\sum_{j,k=1}^{d_{\pi}}e_{j,k}\otimes(\pi_{k,j})^{*}\in C(G)_{2}(\pi)$,
@@
Therefore, $c\in A_{2}(\pi)$, and $c^{*}c\in\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$. Furthermore,
@@
Using $(1-g-\delta)_{+}\precsim_{A^{\alpha}}g$, by part (1) of Lemma 2.5, we get $\delta_{2}>0$ such that $(1-g-2\delta)_{+}\precsim_{A^{\alpha}}(g-\delta_{2})_{+}$. Then, by part (8) of Lemma 2.5, we have $v\in A^{\alpha}$ such that $(1-g-2\delta)_{+}=v^{*}gv$. Clearly, $1_{B(\mathcal{H}_{\pi})}\otimes v\in(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$. Since $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ is a two-sided ideal of $(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$, it follows that
@@
$\|z-1_{B(\mathcal{H}_{\pi})}\otimes 1\|<2(d_{\pi}^{3}\delta+\delta)<1.$
Hence $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ contains the invertible element $z$. ∎
@@
Since $A\rtimes_{\alpha}G$ is simple, $A^{\alpha}$ is a full corner of $A\rtimes_{\alpha}G$ (see [44, Corollary]), hence they are strongly Morita equivalent. Furthermore, since both $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are separable, they are stably isomorphic. ∎
@@ 5. Permanence properties
    $\psi$ is an $(S,F,\varepsilon)$-approximately equivariant central strong order zero map.
@@
We may assume there exist $c_{1},c_{2},\dots,c_{m}\in A$ and $e_{1},e_{2},\dots,e_{m}\in B$ such that $F=\{c_{i}\otimes e_{i}\colon i=1,2,\dots,m\}$ and $\|c_{i}\|\leq 1,\|e_{i}\|\leq 1$ for $i=1,2,\dots,m$. Choose $\delta$ such that $0<\delta<\frac{\varepsilon}{3}$ and $(1-5\delta-\delta^{2})>1-\varepsilon$. Also, choose $\delta_{1}$ such that $\frac{1}{2}<\delta_{1}<1$ and $(1-5\delta-\delta^{2})\delta_{1}^{2}>1-\varepsilon$. Set $z=(x-\delta_{1})_{+}$. Since $\beta$ is trivial, $(A\otimes_{min}B)^{\alpha\otimes\beta}\cong A^{\alpha}\otimes_{min}B^{\beta}$. By Kirchberg’s Slice Lemma (see Lemma 2.8), there exist $x_{1}\in A_{+}$ and $x_{2}\in B_{+}$ such that $x_{1}\otimes x_{2}\precsim_{A\otimes_{min}B}z$ with $\|x_{1}\|=\|x_{2}\|=1$, and $y_{1}\in(A^{\alpha})_{+}$ and $y_{2}\in(B^{\beta})_{+}$ such that $y_{1}\otimes y_{2}\precsim_{(A\otimes_{min}B)^{\alpha\otimes\beta}}y$ with $\|y_{1}\|=\|y_{2}\|=1$. By [31, Lemma 2.2, Lemma 2.4(i)], there exists $w\in A\otimes_{min}B$ such that $\|wxw^{*}-x_{1}\otimes x_{2}\|<\delta^{2}$ and $\|w\|\leq\delta_{1}^{-\frac{1}{2}}<\delta_{1}^{-1}$. Thus there exists $v\in A\otimes_{min}B$, where $v=\sum_{i=1}^{k}v_{i}\otimes w_{i}$ for some $v_{i}\in A$ and $w_{i}\in B$, $i=1,2,\dots,k$ such that
@@
Set $E=\{c_{i}\colon i=1,2,\dots,m\}\cup\{v_{i}\colon i=1,2,\dots,k\}$. Since $B$ and $B^{\beta}$ are simple and unital, by [30, Proposition 2.7(v)], there exist $n,m\in\mathbb{N}$ such that
@@
    $\varphi$ is an $(S,E,\eta)$-approximately equivariant central strong order zero map.
@@
Assuming (1), by [36, Lemma 4.12], we have $b\in A$ such that $1=b^{*}ab\precsim_{A}a$.
For the converse, assuming (2), by parts (1) and (8) of Lemma 2.5, we have $b\in A$ such that $1=b^{*}ab$. By [36, Lemma 4.12], AA is purely infinite. ∎
@@
Since $A$ is simple and purely infinite, by [36, Lemma 4.12], there exists $x\in A$ such that $x^{*}a_{1}x=1$. Therefore, by Theorem 2.16 and Corollary 2.18, we have
@@
By Theorem 5.4, the algebra $A^{\alpha}$ is purely infinite. Since $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are stably isomorphic (see Corollary 4.8), by [30, Theorem 4.23], $A\rtimes_{\alpha}G$ is also purely infinite. ∎
@@
[27, Definition 2.1] Let $A$ be a unital C*-algebra with $A\ncong\mathbb{C}$. Then $A$ is tracially $\mathcal{Z}$-absorbing (or tracially $\mathcal{Z}$-stable) if and only if for any finite set $F\subseteq A$, $\varepsilon>0$, $x\in A_{+}\setminus\{0\}$, and $n\in\mathbb{N}$, there exists a completely positive contractive order zero map $\psi\colon M_{n}\to A$ such that
@@
[36, Proposition 4.16] Let $n\in\mathbb{N}$ and let $\varepsilon>0$. Then there exists $\delta>0$ satisfying the following property. Let $D$ be a C*-algebra and for $j,k=1,2,\dots,n$ let $y_{j,k}\in D$, satisfying (in the unitization of $D$)
@@
By [36, Lemma 4.17], we obtain a completely positive contractive order zero map $\psi\colon M_{n}\to\overline{d_{n_{0}}A^{\alpha}d_{n_{0}}}$ such that
@@
The algebra $A$ is $\mathcal{Z}$-stable, hence by [27, Proposition 2.2] it is tracially $\mathcal{Z}$-stable. Therefore, by Theorem 5.9, $A^{\alpha}$ is tracially $\mathcal{Z}$-stable. $A$ is amenable, so $A^{\alpha}$ is also amenable; thus by [27, Theorem 4.1], $A^{\alpha}$ is $\mathcal{Z}$-stable. By [49, Corollary 3.2], $\mathcal{Z}$-stability is preserved under stable isomorphism, and $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are stably isomorphic (see Corollary 4.8), so the crossed product $A\rtimes_{\alpha}G$ is also $\mathcal{Z}$-stable. ∎
@@
Let $A$ be a unital simple separable infinite-dimensional stably finite C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of a second-countable compact group with the weak tracial Rokhlin property with comparison. Let $\iota\colon A^{\alpha}\to A$ be the inclusion map. Then the map $\mathrm{Cu}(\iota)\colon\mathrm{Cu}(A^{\alpha})\to\mathrm{Cu}(A)$ induces an injective map of ordered semigroups from $\mathrm{Cu}(A^{\alpha})_{+}\cup\{0\}$ to $\mathrm{Cu}(A)_{+}\cup\{0\}$. If, additionally, $A$ has stable rank one, then the map $\mathrm{W}(\iota)\colon\mathrm{W}(A^{\alpha})\to\mathrm{W}(A)$ induces an injective map of ordered semigroups from $\mathrm{W}(A^{\alpha})_{+}\cup\{0\}$ to $\mathrm{W}(A)_{+}\cup\{0\}$.
@@
The proof is essentially the same as that of [36, Proposition 4.22] when we use Proposition 5.12 in place of [36, Proposition 4.21], so we omit it. ∎
@@
The proof is essentially the same as that of [6, Theorem 4.1] when we use Proposition 5.12 in place of [6, Lemma 3.7], so we omit it. ∎
@@
The proof is essentially the same as that of [46, Theorem 4.4], so we omit it. ∎
@@
The proof is essentially the same as that of [46, Theorem 4.6], so we omit it. ∎
@@ 6. The weak tracial Rokhlin property
Similar to condition (4) in [41, Definition 4.1], since we consider the transfer of properties from $A$ to $A^{\alpha}$, the Cuntz comparison condition in $A^{\alpha}$ is necessary to handle certain technical details. Thus, the condition $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}y$ appears essential; without it, we may not be able to complete most of the previous proofs.
@@
Let $\varepsilon>0$, let $F\subseteq A$ be a finite set, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Since $A^{\alpha}$ is simple and non-type I, by [41, Lemma 2.1], there exists a positive element $z\in\overline{yA^{\alpha}y}$ such that 0 is a limit point of $\mathrm{sp}(z)$. Define continuous functions $h,h_{0}\colon[0,1]\rightarrow[0,1]$ by
@@
Then we have $\|x-h_{0}(x)\|\leq\frac{\varepsilon}{2}$. Moreover, since $\|x\|=1$, we have $h(x)\neq 0$. By [41, Lemma 2.6], we can choose a nonzero positive element $x_{0}\in\overline{h(x)Ah(x)}$ such that $x_{0}\precsim_{A}z$. We may assume $\|x_{0}\|=1$.
@@
By [6, Lemma 3.7], $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}z$. Since $z\in\overline{yA^{\alpha}y}$, we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}y$, which is condition (6).
@@
Let $\varepsilon>0$, let $F\subseteq A$ be a finite set, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Since $A^{\alpha}$ is simple and non-type I, by [41, Lemma 2.1], there exists a positive element $z\in\overline{yA^{\alpha}y}$ such that 0 is a limit point of $\mathrm{sp}(z)$.
@@
Since $A^{\alpha}$ is simple and non-type I, by Lemma 2.6, there exist $n+1$ mutually orthogonal, Cuntz equivalent nonzero positive elements in $A^{\alpha}$. Let $e$ be one of them. Thus, by [41, Lemma 2.1], there exists $r\in(\overline{eA^{\alpha}e})_{+}\setminus\{0\}$ such that 0 is a limit point of $\mathrm{sp}(r)$. Define a continuous function $h\colon[0,1]\to[0,1]$ by
@@
Apply [41, Lemma 2.6] to find $x_{0}\in\overline{h(x)Ah(x)}$ such that
@@
By Proposition 6.3, we only need to show that $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}d$. Since $(1-d-\varepsilon)_{+}\precsim_{A}x_{0}\precsim_{A}r$, by [6, Lemma 3.7], we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}r$. Thus $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}e$. Let $\tau\in\mathrm{QT}(A^{\alpha})$. Then $d_{\tau}((1-d-\varepsilon)_{+})\leq d_{\tau}(e)$, and by Lemma 2.7,
@@
Using [6, Theorem 4.1] in the first step and (6.1) in the second step, we have $\mathrm{rc}(A^{\alpha})\leq\mathrm{rc}(A)<1-\frac{2}{n}$. Since (6.2) holds for all $\tau\in\mathrm{QT}(A^{\alpha})$, we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}d$. ∎
@@
Let $A$ be a unital, simple separable infinite-dimensional stably finite C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of a second-countable compact group with the weak tracial Rokhlin property. Let $\iota\colon A^{\alpha}\to A$ be the inclusion map. Then the map $\mathrm{Cu}(\iota)\colon\mathrm{Cu}(A^{\alpha})\to\mathrm{Cu}(A)$ induces an injective map of ordered semigroups from $\mathrm{Cu}(A^{\alpha})_{+}\cup\{0\}$ to $\mathrm{Cu}(A)_{+}\cup\{0\}$. If, additionally, $A$ has stable rank one, then the map $\mathrm{W}(\iota)\colon\mathrm{W}(A^{\alpha})\to\mathrm{W}(A)$ induces an injective map of ordered semigroups from $\mathrm{W}(A^{\alpha})_{+}\cup\{0\}$ to $\mathrm{W}(A)_{+}\cup\{0\}$.

### v2

@@ 4. Simplicity of the fixed point algebra and the crossed product
Since $C(G)$ is amenable and separable, by the Choi–Effros lifting theorem, for $\delta$, $F=\{1_{A}\}$, and $S=\{\pi_{j,k},(\pi_{j,k})^{*}\colon j,k=1,2,\dots,d_{\pi}\}$, we obtain a contractive positive element $g\in A^{\alpha}$ such that $(1-g-\delta)_{+}\precsim_{A^{\alpha}}g$ and an $(S,F,\delta)$-approximately equivariant central strongly order zero completely positive contractive map $\psi\colon C(G)\to\overline{gAg}$ with $\psi(1)=g^{\frac{1}{2}}$ such that for all $f_{1},f_{2}\in S$, $\|g^{\frac{1}{2}}\psi(f_{1}f_{2})-\psi(f_{1})\psi(f_{2})\|<\delta$. As in Lemma 3.6, we will assume that $\psi$ is equivariant. Then
$\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\psi\colon B(\mathcal{H}_{\pi})\otimes C(G)\to\overline{(1_{B(\mathcal{H}_{\pi})}\otimes g)(B(\mathcal{H}_{\pi})\otimes A)(1_{B(\mathcal{H}_{\pi})}\otimes g)}.$
is an equivariant completely positive contractive map, and $(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\psi)(1)=1_{B(\mathcal{H}_{\pi})}\otimes g^{\frac{1}{2}}$.
Let $(e_{j,k})_{j,k=1,2,\dots,d_{\pi}}$ be the standard matrix units of $B(\mathcal{H}_{\pi})$. Let
@@
Since $\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\psi$ is equivariant and $\pi^{*}\in C(G)_{2}(\pi)$, a simple calculation shows that $c\in A_{2}(\pi)$. Indeed, for $g\in G$, since $\pi^{*}=\sum_{j,k=1}^{d_{\pi}}e_{j,k}\otimes(\pi_{k,j})^{*}\in C(G)_{2}(\pi)$, we have
@@
Therefore, $c\in A_{2}(\pi)$, and $c^{*}c\in\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$. Furthermore, since $\pi\pi^{*}=1$, we have
@@
Using $(1-g-\delta)_{+}\precsim_{A^{\alpha}}g$, by (1) of Lemma 2.5, we obtain $\delta_{2}>0$ such that $(1-g-2\delta)_{+}\precsim_{A^{\alpha}}(g-\delta_{2})_{+}$. Then, by (8) of Lemma 2.5, we have $v\in A^{\alpha}$ such that $(1-g-2\delta)_{+}=v^{*}gv$. Clearly, $1_{B(\mathcal{H}_{\pi})}\otimes v\in(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$. Since $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ is a two-sided ideal of $(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$, it follows that
@@
$\|z-1_{B(\mathcal{H}_{\pi})}\otimes 1\|<2(d_{\pi}^{3}\delta+\delta)\leq 1.$
Hence, $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ contains the invertible element $z$. ∎
@@
Since $A\rtimes_{\alpha}G$ is simple, $A^{\alpha}$ is a full corner of $A\rtimes_{\alpha}G$ (see [46, Corollary]), hence they are strongly Morita equivalent. Furthermore, since both $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are separable, they are stably isomorphic. ∎
@@ 5. Permanence properties
The following proposition tells us that the compact group action with the weak tracial Rokhlin property of comparison is preserved under taking tensor products. However, for the technical handling of Cuntz comparison, it seems we can only require the action of one group to be trivial. Furthermore, since we need to use Kirchberg’s Slice Lemma, we employ the minimal tensor product.
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately equivariant central strongly order zero map.
@@
We may assume there exist $c_{1},c_{2},\dots,c_{m}\in A$ and $e_{1},e_{2},\dots,e_{m}\in B$ such that $F=\{c_{i}\otimes e_{i}\colon i=1,2,\dots,m\}$ and $\|c_{i}\|\leq 1,\|e_{i}\|\leq 1$ for $i=1,2,\dots,m$. Choose $\delta$ such that $0<\delta<\frac{\varepsilon}{3}$ and $(1-5\delta-\delta^{2})>1-\varepsilon$. Also, choose $\delta_{1}$ such that $\frac{1}{2}<\delta_{1}<1$ and $(1-5\delta-\delta^{2})\delta_{1}^{2}>1-\varepsilon$. Set $z=(x-\delta_{1})_{+}$. Since $\beta$ is trivial, $(A\otimes_{min}B)^{\alpha\otimes\beta}\cong A^{\alpha}\otimes_{min}B^{\beta}$. By Kirchberg’s Slice Lemma (see Lemma 2.8), there exist $x_{1}\in A_{+}$ and $x_{2}\in B_{+}$ such that $x_{1}\otimes x_{2}\precsim_{A\otimes_{min}B}z$ with $\|x_{1}\|=\|x_{2}\|=1$, and $y_{1}\in(A^{\alpha})_{+}$ and $y_{2}\in(B^{\beta})_{+}$ such that $y_{1}\otimes y_{2}\precsim_{(A\otimes_{min}B)^{\alpha\otimes\beta}}y$ with $\|y_{1}\|=\|y_{2}\|=1$. By [32, Lemma 2.2, Lemma 2.4(i)], there exists $w\in A\otimes_{min}B$ such that $\|wxw^{*}-x_{1}\otimes x_{2}\|<\delta^{2}$ and $\|w\|\leq\delta_{1}^{-\frac{1}{2}}<\delta_{1}^{-1}$. Thus there exists $v\in A\otimes_{min}B$, where $v=\sum_{i=1}^{k}v_{i}\otimes w_{i}$ for some $v_{i}\in A$ and $w_{i}\in B$, $i=1,2,\dots,k$ such that
@@
Set $E=\{c_{i}\colon i=1,2,\dots,m\}\cup\{v_{i}\colon i=1,2,\dots,k\}$. Since $B$ and $B^{\beta}$ are simple and unital, by [31, Proposition 2.7(v)], there exist $n,m\in\mathbb{N}$ such that
@@
    $\varphi$ is an $(S,E,\eta)$-approximately equivariant central strongly order zero map.
@@
Now we turn to pure infiniteness. For convenience, we state the following lemma separately.
@@
Assuming (1), by [37, Lemma 4.12], we have $b\in A$ such that $1=b^{*}ab\precsim_{A}a$.
For the converse, assuming (2), by parts (1) and (8) of Lemma 2.5, we have $b\in A$ such that $1=b^{*}ab$. By [37, Lemma 4.12], AA is purely infinite. ∎
@@
Since $A$ is simple and purely infinite, by [37, Lemma 4.12], there exists $x\in A$ such that $x^{*}a_{1}x=1$. Therefore, by Theorem 2.16 and Corollary 2.18, we have
@@
By Theorem 5.4, the algebra $A^{\alpha}$ is purely infinite. Since $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are stably isomorphic (see Corollary 4.9), by [31, Theorem 4.23], $A\rtimes_{\alpha}G$ is also purely infinite. ∎
@@
[28, Definition 2.1] Let $A$ be a unital C*-algebra with $A\ncong\mathbb{C}$. Then $A$ is tracially $\mathcal{Z}$-absorbing (or tracially $\mathcal{Z}$-stable) if and only if for any finite set $F\subseteq A$, $\varepsilon>0$, $x\in A_{+}\setminus\{0\}$, and $n\in\mathbb{N}$, there exists a completely positive contractive order zero map $\psi\colon M_{n}\to A$ such that
@@
[37, Proposition 4.16] Let $n\in\mathbb{N}$ and let $\varepsilon>0$. Then there exists $\delta>0$ satisfying the following property. Let $D$ be a C*-algebra and for $j,k=1,2,\dots,n$ let $y_{j,k}\in D$, satisfying (in the unitization of $D$)
@@
By [37, Lemma 4.17], we obtain a completely positive contractive order zero map $\psi\colon M_{n}\to\overline{d_{n_{0}}A^{\alpha}d_{n_{0}}}$ such that
@@
The algebra $A$ is $\mathcal{Z}$-stable, hence by [28, Proposition 2.2] it is tracially $\mathcal{Z}$-stable. Therefore, by Theorem 5.9, $A^{\alpha}$ is tracially $\mathcal{Z}$-stable. $A$ is amenable, so $A^{\alpha}$ is also amenable; thus by [28, Theorem 4.1], $A^{\alpha}$ is $\mathcal{Z}$-stable. By [51, Corollary 3.2], $\mathcal{Z}$-stability is preserved under stable isomorphism, and $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are stably isomorphic (see Corollary 4.9), so the crossed product $A\rtimes_{\alpha}G$ is also $\mathcal{Z}$-stable. ∎
@@
The following proposition is a generalization of [37, Proposition 4.22]. This result is weaker than its counterpart for compact group actions with the Rokhlin property; namely, we cannot obtain an embedding of the Cuntz semigroup of the fixed point algebra into the Cuntz semigroup of the original algebra, but only an embedding between the subsemigroups consisting of purely positive elements (i.e., elements not Cuntz equivalent to a projection).
@@
Let $A$ be a unital simple separable infinite-dimensional stably finite C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of a second-countable compact group with the weak tracial Rokhlin property with comparison. Let $\iota\colon A^{\alpha}\to A$ be the inclusion map. Then
  1. (1)
    the map $\mathrm{W}(\iota)\colon\mathrm{W}(A^{\alpha})\to\mathrm{W}(A)$ induces an injective map of ordered semigroups from $\mathrm{W}(A^{\alpha})_{+}\cup\{0\}$ to $\mathrm{W}(A)_{+}\cup\{0\}$;
  2. (2)
    the map $\mathrm{Cu}(\iota)\colon\mathrm{Cu}(A^{\alpha})\to\mathrm{Cu}(A)$ induces an injective map of ordered semigroups from $\mathrm{Cu}(A^{\alpha})_{+}\cup\{0\}$ to $\mathrm{Cu}(A)_{+}\cup\{0\}$.
@@
The proof is essentially the same as that of [37, Proposition 4.22] when we use Proposition 5.12 in place of [37, Proposition 4.21], so we omit it. ∎
The following proposition establishes a correspondence between the 2-quasitraces on the fixed point algebra under the action of a second-countable compact group with the weak tracial Rokhlin property of comparison and the 2-quasitraces on the original algebra.
@@
The proof is essentially the same as that of [7, Theorem 4.1] when we use Proposition 5.12 in place of [7, Lemma 3.7], so we omit it. ∎
@@
The proof is essentially the same as that of [48, Theorem 4.4], so we omit it. ∎
@@
The proof is essentially the same as that of [48, Theorem 4.6], so we omit it. ∎
@@ 6. The weak tracial Rokhlin property
    $\psi$ is an $(S,F,\varepsilon)$-approximately central equivariant strongly order zero map.
@@
Similar to condition (4) in [43, Definition 4.1], since we consider the transfer of properties from $A$ to $A^{\alpha}$, the Cuntz comparison condition in $A^{\alpha}$ is necessary to handle certain technical details. Thus, the condition $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}y$ appears essential; without it, we may not be able to complete most of the previous proofs.
@@
Let $\varepsilon>0$, let $F\subseteq A$ be a finite set, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Since $A^{\alpha}$ is simple and non-type I, by [43, Lemma 2.1], there exists a positive element $z\in\overline{yA^{\alpha}y}$ such that 0 is a limit point of $\mathrm{sp}(z)$. Define continuous functions $h,h_{0}\colon[0,1]\rightarrow[0,1]$ by
@@
Then we have $\|x-h_{0}(x)\|\leq\frac{\varepsilon}{2}$. Moreover, since $\|x\|=1$, we have $h(x)\neq 0$. By [43, Lemma 2.6], we can choose a nonzero positive element $x_{0}\in\overline{h(x)Ah(x)}$ such that $x_{0}\precsim_{A}z$. We may assume $\|x_{0}\|=1$.
@@
By [7, Lemma 3.7], $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}z$. Since $z\in\overline{yA^{\alpha}y}$, we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}y$, which is condition (6).
@@
Let $\varepsilon>0$, let $F\subseteq A$ be a finite set, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Since $A^{\alpha}$ is simple and non-type I, by [43, Lemma 2.1], there exists a positive element $z\in\overline{yA^{\alpha}y}$ such that 0 is a limit point of $\mathrm{sp}(z)$.
@@
Since $A^{\alpha}$ is simple and non-type I, by Lemma 2.6, there exist $n+1$ mutually orthogonal, Cuntz equivalent nonzero positive elements in $A^{\alpha}$. Let $e$ be one of them. Thus, by [43, Lemma 2.1], there exists $r\in(\overline{eA^{\alpha}e})_{+}\setminus\{0\}$ such that 0 is a limit point of $\mathrm{sp}(r)$. Define a continuous function $h\colon[0,1]\to[0,1]$ by
@@
Apply [43, Lemma 2.6] to find $x_{0}\in\overline{h(x)Ah(x)}$ such that
@@
By Proposition 6.3, we only need to show that $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}d$. Since $(1-d-\varepsilon)_{+}\precsim_{A}x_{0}\precsim_{A}r$, by [7, Lemma 3.7], we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}r$. Thus $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}e$. Let $\tau\in\mathrm{QT}(A^{\alpha})$. Then $d_{\tau}((1-d-\varepsilon)_{+})\leq d_{\tau}(e)$, and by Lemma 2.7,
@@
Using [7, Theorem 4.1] in the first step and (6.1) in the second step, we have $\mathrm{rc}(A^{\alpha})\leq\mathrm{rc}(A)<1-\frac{2}{n}$. Since (6.2) holds for all $\tau\in\mathrm{QT}(A^{\alpha})$, we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}d$. ∎
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central equivariant strongly order zero map.
@@
Let $A$ be a unital simple separable infinite-dimensional stably finite C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of a second-countable compact group with the weak tracial Rokhlin property. Let $\iota\colon A^{\alpha}\to A$ be the inclusion map. Then
  1. (1)
    the map $\mathrm{W}(\iota)\colon\mathrm{W}(A^{\alpha})\to\mathrm{W}(A)$ induces an injective map of ordered semigroups from $\mathrm{W}(A^{\alpha})_{+}\cup\{0\}$ to $\mathrm{W}(A)_{+}\cup\{0\}$;
  2. (2)
    the map $\mathrm{Cu}(\iota)\colon\mathrm{Cu}(A^{\alpha})\to\mathrm{Cu}(A)$ induces an injective map of ordered semigroups from $\mathrm{Cu}(A^{\alpha})_{+}\cup\{0\}$ to $\mathrm{Cu}(A)_{+}\cup\{0\}$.